# Characterizing the Sensitivity to Individual Bit Flips in Client-Side Operations of the CKKS Scheme

### v1

@@ II-C CKKS’ Insights
To better understand the parameter selection in CKKS, it is helpful to clarify the concept of precision. In the case of OpenFHE with real-valued inputs only, the scaling parameter $\Delta$ determines how many bits are used to encode the fractional part of the input values, while the difference between the initial modulus $q_{0}$ and $\Delta$ indicates the number of bits allocated to the integer part.
For example, using $q_{0}=60$ bits y $\Delta=50$ bits allows the encrypted representation of real numbers with approximately 10 bits of precision for the integer part and 50 bits for the fractional part.
@@
Encryption in the CKKS scheme requires either a secret key $sk$ or a public key $pk$. The secret key is generated by sampling a vector $s$ of length $N$ generally from a ternary distribution with Hamming weight $h$, and then setting $sk=(1,s)$.
The public key is constructed by sampling $a_{0}$ from $\mathbb{Z}_{Q}$, i.e., from the set of positive integers less than the modulus $Q$, and an error term $e_{0}$ from a discrete Gaussian distribution with standard deviation $\sigma=3.2$. The public key is then defined as:
@@
where the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
@@
The key idea behind encoding in CKKS is to transform the input vector into a polynomial. Different libraries implementing CKKS vary in how they perform this transformation using IFFT/FFT-based methods. In all cases, the standard cyclic transform must be adapted to a negacyclic version to ensure operations remain within the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$ used by the scheme.
The transforms employed by HEAAN 1.0 and OpenFHE 1.2.3 are negacyclic and include scheme-specific optimizations, commonly referred to as specialized FFT/IFFT. These versions incorporate several improvements originally proposed by Bernstein [19, 20]. One of their main advantages is that they allow the transform to be applied directly to vectors of size $\leq N/2$, without expanding them to size $N$, which significantly reduces computational cost.
Given a real-valued input vector of size $n\leq N/2$, the encoding process begins by padding the vector with zeros to reach the next power of two, denoted $n^{\prime}$. A specialized inverse FFT is then applied, producing a complex vector of length $n^{\prime}$. From this, a polynomial of degree $N-1$ is constructed.
First, the gap parameters $=\frac{N/2}{n^{\prime}}$ is computed, and the lower half of the polynomial (coefficients of degree 0 to $N/2-1$) is filled with the real parts of the transformed vector, with assignments spaced at intervals of size gap. Unassigned coefficients are set to zero.
Finally, as is illustrated in Fig. 3, each coefficient is multiplied by $\Delta$, and then rounded to the nearest integer using the coordinate-wise randomized rounding method described in [21].
Figure 3: Encoding scheme in CKKS using HEAAN/OpenFHE. The example shows an input vector with $n=2$ values, represented in a polynomial ring of size $N=8$, where the gap between consecutive coefficients is 2.
@@
where $a$ is a polynomial sampled from a uniform distribution, $e$ is a noise term sampled from a discrete Gaussian distribution, and the notation $[\cdots]_{Q}$ denotes reduction modulo $Q$.
An alternative way is encrypting by using the public key, $pk=(p_{0},p_{1})$. In this case, encryption is defined as:
@@
Decryption in both cases (secret key and public key encryption) uses the secret key $sk=(1,s)$ and the encrypted text $(c_{0},c_{1})$ as follows:
$m^{\prime}=[c_{0}+c_{1}\times s]_{Q}$ (4)
@@
To decode a plaintext polynomial $m$ with $N$ coefficients, one must reconstruct a complex-valued vector of $n^{\prime}$ elements by reversing the encoding process illustrated in Fig. 3. The $i$-th element ($i\leq N/2$) of the resulting vector is formed by using the real part of the $i$-th coefficient of the plaintext, and the imaginary part from the coefficient at position $i+N/2$.
This complex vector is then divided by $\Delta$, and a specialized FFT is applied to recover the approximate original input.
Due to rounding effects during encoding and decoding, the final result may contain small approximation errors. Corollary 3 in [22] shows that the worst-case distance between the input vector $z$ and the output vector $z^{\prime}$ (after encoding and decoding) at each position is approximately: $N/(\pi\times\Delta)$.
[…]
Security Against Secret Key Attacks
To defend against secret key recovery attacks, OpenFHE performs additional operations before applying the specialized FFT during the decoding process. Specifically, a Gaussian error is injected prior to this step, with its standard deviation determined by the vector being decoded. This mechanism is critical: in our bit-flip experiments–when security analysis is not disabled–the software throws an exception during this stage, resulting in premature termination of execution.
This behavior demonstrates the role of these protective measures in preventing potential side-channel or fault-injection attacks that could compromise the secret key.
II-D RNS and NTT Optimizations
In the full RNS/NTT variant of CKKS [3], the Residue Number System (RNS) is used to split large polynomial coefficients into multiple smaller values, enabling more efficient computations. This approach allows arithmetic in $\mathbb{Z}_{Q}$ to be performed using standard machine word sizes, by working with small moduli that simplify computation.
The magnitude of the modulus $Q$ can range from hundreds to thousands of bits, depending on the use case. However, modern computing systems typically operate with word sizes of 64 bits or less, making direct operations on such large coefficients impractical. To address this, polynomials (keys, plaintexts, ciphertexts) are represented using an equivalent decomposition in the RNS form. Each coefficient is represented modulo a small prime $q_{i}$, to generate $Q$ like Eq. 1
This decomposition enables addition and multiplication in $\mathbb{Z}_{Q}$ using standard machine word operations. All other parameters, as well as the encoding and decoding procedures, remain consistent with the vanilla CKKS variant.
Polynomial multiplications are performed efficiently using the Number Theoretic Transform (NTT), which reduces the computational complexity of the operation from $\mathcal{O}(n^{2})$ to $\mathcal{O}(n\log n)$. Addition of polynomials and multiplication by a scalar are both $\mathcal{O}(n)$ operations in both the coefficient and evaluation (NTT) representations.
By default, the CKKS scheme assumes that polynomials are represented in the NTT domain (evaluation representation). When certain operations require polynomials in coefficient form, an inverse NTT (iNTT) is applied to transform ciphertexts from the evaluation domain back to the coefficient domain. However, within the scope of this work, as all computations are performed on the client side, this change of representation does not occur.
The procedures for encoding, encryption, decryption, and decoding remain virtually unchanged, except that all operations are now performed using the RNS and NTT representations.
@@ III Error Resilience Analysis
This section presents a theoretical analysis of the error induced by a single bit-flip at different stages of the CKKS scheme. We address both the vanilla and the full RNS variant versions. Our aim is:
  • Deriving analytical bounds in the case without RNS or NTT.
  • Identifying the most sensitive coefficients.
  • Providing an intuitive description of error behavior in practical implementations.
Although a closed-form analytical bound can not be derived for the full RNS variant version due to the complexity of modular representations, a qualitative analysis of the observed patterns is provided.
III-A Error Model for a Single Bit-Flip
We adopt a single bit-flip model, where a single bit is altered. In practice, such faults may be caused by electromagnetic interference, hardware failures, or cosmic rays [12, 15, 23, 13, 24, 17].
@@ III-A Error Model for a Single Bit-Flip
We model this fault as an error vector $\boldsymbol{e}_{i,j}\in\mathbb{R}^{N}$ whose only non-zero component corresponds to bit $j$ of coefficient $i$:
$\mathbf{e}_{i,j}=(0,\dots,0,e_{j},0,\dots,0)$ (6)
This model enables us to track how a single bit-flip propagates through the cryptographic pipeline.
@@ III-B CKKS Encoding via DFT
As discussed in Section II, one of the key components of encoding and decoding in CKKS is FFT. Although CKKS uses a specialized inverse FFT for computational efficiency, this transformation is typically implemented using optimized algorithms that obscure its underlying linear structure.
Instead, we rely on the negacyclic Discrete Fourier Transform (DFT)–an algebraically equivalent transformation explicitly expressed as matrix multiplication. This formulation allows us to:
@@
While this choice sacrifices performance, it offers significant pedagogical advantages, allowing for an intuitive understanding of how bit-flip errors are amplified or distributed during encoding.
@@ III-B1 Encoding via Negacyclic DFT
In order to use the DFT and its inverse for encoding to obtain results equivalent to those produced by implementations such as HEAAN and OpenFHE, some preprocessing is required: the input vector must be resized to length $N$ and reordered to satisfy the requirements of those systems. However, these adjustments do not affect the behavior of the error model under analysis and will be omitted for simplicity.
@@
Finally, the output undergoes the same reordering process as in encoding, and the $n$ components of $z^{\prime}$ are extracted. Since the DFT output is complex-valued while OpenFHE supports only real-valued encryption, the real part is extracted to obtain the final approximation of z
III-C Theoretical Estimation of Encoding Error in Vanilla CKKS
This section focuses on the impact of single-bit errors during the encoding phase of the CKKS scheme.
Let us consider the case in which an error $\boldsymbol{e}_{i,j}$ is introduced by flipping a single bit in coefficient $i$ of the plaintext vector $\boldsymbol{m}$, specifically at the $j$-th bit position. The resulting perturbed plaintext is $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$, and decoding proceeds as in Eq. 9:
@@ III-C Theoretical Estimation of Encoding Error in Vanilla CKKS
Since a single bit-flip affects only one coefficient, the resulting error is localized. The difference between the decoded vectors $\boldsymbol{z^{\prime}}$ and $\boldsymbol{z^{\prime\prime}}$ is given by:
@@ III-C1 Bit-Flip Injection during Encoding
To better understand the analytical expression of the error in Eq. (10), we simulate flipping each individual bit in the plaintext coefficients. Let $j\in\{0,\dots,63\}$ denotes the bit position in a 64-bit unsigned integer. For each iteration, we flip the $j$-th bit of a coefficient, introducing an error with integer magnitude $e_{j}=2^{j}$. This represents the maximum jump in value induced by flipping bit $j$ in a coefficient, which is stored as an unsigned integer–as in the case of plaintext coefficients.
We repeat this process for all coefficients, grouping the results into $N=4$ blocks of 64 iterations each, and use $\Delta=2^{50}$. Let $\boldsymbol{z}^{\prime}$ be the decoded vector without errors, and $\boldsymbol{z}^{\prime\prime}$ the result after applying the bit-flip and decoding. Using Eq. (10), we define:
@@
where $\mathbf{e}_{i,j}$ is a zero vector of length $N$ with value $e_{j}$ at position $i$.
We also take the real part of the vector before computing the L2 norm, to simulate the encoding of real-valued data, as in OpenFHE, in contrast to HEAAN, which allows complex inputs. This step is important not only to maintain consistency with the backend used in later sections, but also because it exposes a specific behavior that would otherwise not manifest when keeping the complex part.
Figure 4: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding. The horizontal axis shows the bit-flip position $i*64+j$, and the vertical axis shows $L_{2}(i,j)$. Ring dimension $N=4$; scaling factor $\Delta=2^{50}$.
@@ III-C2 Effect of the Scaling Factor on Bit-Flip Error
Figure 5: $L_{2}$ norm of the error after a single bit-flip in the plaintext during encoding, for various scaling factors $\Delta$. The horizontal axis indicates the bit position $j$; the vertical axis is in semi-log scale. Ring dimension: $N=4$.
As shown in Fig. 5, increasing $\Delta$ significantly reduces the error magnitude and increases the number of resilient bits. This occurs because the impact of errors becomes noticeable only when the flipped bit corresponds to or exceeds the magnitude of $\Delta$. In other words, the larger the scaling factor, the more initial bits remain unaffected, yielding smaller error norms and higher bit-level fault tolerance.
This empirical observation is analytically confirmed by equation 11, which shows that the error is inversely proportional to the scaling factor. When an error $e_{i,j}$ is divided by $\Delta$, noise components with magnitudes smaller than the bit precision of $\Delta$ are effectively suppressed in the final result. This mathematical relationship explains why larger scaling factors provide enhanced fault resilience: they act as a natural filter that attenuates the impact of small-magnitude bit flips on the decoded output.
However, this resilience comes at the cost of reduced fractional precision in the encoded message (see Section II). The selection of $\Delta$ therefore requires careful consideration of the trade-off between error tolerance and numerical precision, balancing the competing demands of fault resilience and computational accuracy.
@@ III-D Theoretical Estimation of Error during Encryption using Vanilla CKKS
This section analyzes three scenarios involving a single bit-flip in vanilla CKKS:
  1. 1.
    A bit-flip in the plaintext (before encryption).
  2. 2.
    A bit-flip in the ciphertext (in one of the ciphertext polynomials).
  3. 3.
    A bit-flip in the plaintext (before decoding).
Under public-key encryption, a plaintext vector $\mathbf{m}$ is encoded into a polynomial $\boldsymbol{m}(X)$, encrypted as in Eq. 3, and decrypted using Eq. 4.
III-D1 Bit‐Flip in the Plaintext Before Encryption
Suppose a bit-flip introduces an error $\boldsymbol{e}_{i,j}$ in the $i$-th coefficient at bit position $j$, so the plaintext becomes:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j},\quad\boldsymbol{e}_{i,j}=e_{j}\,X^{i-1}$ (12)
Encrypting $\boldsymbol{m}^{\prime}$ using the public key (as in Eq. 3) yields:
$\begin{split}ct^{\prime}&=\bigl{(}[\,\boldsymbol{m}^{\prime}+p_{0}v+e_{1}\,]_{Q},\;[\,p_{1}v+e_{2}\,]_{Q}\bigr{)}\\ &=\bigl{(}[\,\boldsymbol{m}+e_{j}\,X^{i-1}+p_{0}v+e_{1}\,]_{Q},\;c_{1}\bigr{)}=\bigl{(}c_{0}^{\prime},\,c_{1}\bigr{)}.\end{split}$ (13)
The error affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Upon decryption (using Eq. 4):
$\boldsymbol{m}^{\prime\prime}=[\,c_{0}^{\prime}+c_{1}\,s\,]_{Q}=\boldsymbol{m}+e_{j}\,X^{i-1}+(e_{1}+e_{2}\,s)$ (14)
However, the error magnitude grows exponentially with the flipped bit position, as each bit represents a successive power of 2 in the binary representation. Consequently, an exhaustive bit-flip analysis across all coefficient positions, with L2 norm computation for each perturbation, yields the exponential error scaling pattern illustrated in Fig. 4 (SectionIII-C). that the bit-flip impact remains localized to the position $i$ and scales linearly with $e_{j}$.
The error induced by the bit-flip affects only the ciphertext component $c_{0}^{\prime}$ in coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Decoding this plaintext will therefore produce an error consistent with that discussed in Section III-C.
III-D2 Bit‐Flip in the Ciphertext
If the bit-flip occurs in the ciphertext, two sub-cases arise:

### v3

@@ II-D Encoding/Encryption/Decryption/Decoding
To better understand the parameter selection in CKKS, it is helpful to clarify the concept of precision. Some CKKS implementations use real-valued inputs only, the scaling parameter $\Delta$ determines how many bits are used to encode the fractional part of the input values, and the difference between the initial modulus $q_{0}$ and $\Delta$ indicates the number of bits allocated to the integer part. For example, using $q_{0}=60$ bits and $\Delta=50$ bits allows the encrypted representation of real numbers with approximately 10 bits of precision for the integer part and 50 bits for the fractional part.
@@
Encryption in the CKKS scheme requires either a secret key $sk$ or a public key $pk$. The secret key is generated as $sk=(1,s)$ by sampling a vector $s$ of length $N$ generally from a ternary distribution with Hamming weight $h$. The public key is constructed by sampling $a_{0}$ from $\mathbb{Z}_{Q}$ (i.e., from the set of positive integers less than the modulus $Q$) and an error term $e_{0}$ from a discrete Gaussian distribution with standard deviation $\sigma=3.2$, as follows:
@@
where the notation $\cdots]_{Q}$ denotes reduction modulo $Q$.
@@
The key idea behind encoding in CKKS is to transform the input vector into a polynomial. Different libraries implementing CKKS vary in how they perform this transformation using IFFT/FFT-based methods. In all cases, the standard cyclic transform must be adapted to a negacyclic version to ensure that operations remain within the ring $\mathbb{Z}_{q}[X]/(X^{N}+1)$ used by the scheme.
The transforms employed by the most adopted CKKS implentations are negacyclic and include scheme-specific optimizations, commonly referred to as specialized FFT/IFFT. These versions incorporate several improvements originally proposed by Bernstein [15, 16]. One of their main advantages is that they allow the transform to be applied directly to vectors of size $\leq N/2$, without expanding them to size $N$, which significantly reduces computational cost.
Given a complex-valued input vector of size $n\leq N/2$, the encoding process begins by padding the vector with zeros to reach the next power of two, denoted $n^{\prime}$. A specialized inverse FFT is then applied, producing a complex vector of length $n^{\prime}$. From this, a polynomial of degree N1N-1 is constructed.
First, the gap parameter $gap=\frac{N/2}{n^{\prime}}$ is computed, and the lower half of the polynomial (coefficients of degree 0 to N/21N/2-1) is filled with the real parts of the transformed vector, with assignments spaced at intervals of size gap. The upper half does the same using the imaginary part of the transformed vector. Unassigned coefficients are set to zero. This process is depicted in Figure 3. As part of this encoding stage, each coefficient is multiplied by $\Delta$, and then rounded to the nearest integer using the coordinate-wise randomized rounding method described in [17].
Figure 3: Standard encoding scheme in CKKS. The example shows an input vector with $n=2$ values, represented in a polynomial ring of size $N=8$, where the gap between consecutive coefficients is 2.
@@
where $a$ is a polynomial sampled from a uniform distribution, $e$ is a noise term sampled from a discrete Gaussian distribution, and the notation $\cdots]_{Q}$ denotes reduction modulo $Q$.
An alternative procedure is to encrypt by using the public key, $pk=(p_{0},p_{1})$. In this case, encryption is defined as:
@@
Decryption in both secret key and public key cases uses the secret key $sk=(1,s)$ and the encrypted text $(c_{0},c_{1})$ as follows:
$m=[c_{0}+c_{1}\times s]_{Q}$ (4)
@@
Decoding a plaintext polynomial $m$ with $N$ coefficients implies reconstructing a complex-valued vector of $n^{\prime}$ elements by reversing the encoding process illustrated in Figure 3. The $i$-th element ($i\leq N/2$) of the resulting vector is formed by using the real part of the $i$-th coefficient of the plaintext, and the imaginary part from the coefficient at position $i+N/2$. This complex vector is then divided by $\Delta$, and a specialized FFT is applied to recover the approximate original input.
Even when simply encoding and decoding an input vector without performing any computation, small approximation errors may arise due to rounding effects. Corollary 3 in [18] shows that the worst-case distance between the input vector $z$ and the output vector $z^{\prime}$ (after encoding and decoding) at each position is approximately: $N/(\pi\times\Delta)$. With a sufficiently large $\Delta$, this difference becomes negligible. Given that a typical ring size $N$ ranges from $2^{12}-2^{16}$, using a scaling factor on the order of $2^{20}$ is enough to ensure a negligible approximation error.
@@ III Error Resilience Analysis
This section presents a theoretical analysis of the error induced by a single bit flip at different stages of the CKKS scheme, in its vanilla and full RNS variants. The goal of this analysis is:
  • Determining analytical error bounds in the vanilla variant (no RNS/NTT).
  • Identifying the most sensitive polynomial coefficients.
  • Enabling a better understanding of error behavior in practical implementations.
Although a closed-form analytical bound can not be derived for the full RNS variant due to the complexity of modular representations, a qualitative analysis of the observed patterns is provided.
III-A Error Model for a Single Bit Flip
We adopt a single bit-flip (soft error) model, where a single bit is altered. In practice, such faults may be caused by electromagnetic interference [7, 10], hardware failures [19, 8], or cosmic rays [20, 12].
@@ III-A Error Model for a Single Bit Flip
We model this fault as an error polynomial $\boldsymbol{e}_{i,j}\in\mathbb{Z}[X]$, whose only non-zero coefficient is the (i1)(i-1)-th, corresponding to an error at bit $j$ of coefficient i1i-1:
$\boldsymbol{e}_{i,j}=e_{j}\,X^{i-1}$ (6)
This model enables us to track how a single bit flip propagates through the cryptographic pipeline.
@@ III-B CKKS Encoding via DFT
As discussed in Section II, the Fast Fourier Transform (FFT) is one of the key components of encoding and decoding in CKKS. Although CKKS uses a specialized inverse FFT for computational efficiency, this transformation is typically implemented using optimized algorithms that obscure its underlying linear structure. Instead, we rely on the negacyclic Discrete Fourier Transform (DFT), an algebraically equivalent transformation explicitly expressed as matrix multiplication. This formulation allows us to:
@@
While this choice sacrifices performance, it offers significant pedagogical advantages, allowing for an intuitive understanding of how bit-flip errors are amplified or distributed during the encoding stage.
@@ III-B1 Encoding via Negacyclic DFT
To use the DFT and its inverse for encoding to obtain results equivalent to those produced by standard implementations, some preprocessing is required: the input vector must be resized to length $N$ and reordered. However, these adjustments do not affect the behavior of the error model under analysis and are omitted for simplicity.
@@
Finally, the output undergoes the same reordering process as in encoding, and the $n$ components of $z^{\prime}$ are extracted. Since the DFT output is complex-valued, the real part is extracted to obtain the final approximation of z
III-C Theoretical Estimation of Encoding Error on vanilla CKKS
This section focuses on the impact of single-bit errors during the encoding phase of the CKKS scheme. Let us consider the case in which an error $\boldsymbol{e}_{i,j}$ is introduced by flipping a single bit in the coefficient $i$ of the plaintext vector $\boldsymbol{m}$, specifically at the $j$-th bit position. The resulting perturbed plaintext is $\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$, and decoding proceeds as in Equation 9:
@@ III-C Theoretical Estimation of Encoding Error on vanilla CKKS
Since a single bit flip affects only one coefficient, the resulting error is localized. The difference between the decoded vectors $\boldsymbol{z^{\prime}}$ and $\boldsymbol{z^{\prime\prime}}$ is given by:
@@ III-C1 Bit-Flip Injection during Encoding
To better understand the analytical expression of the error in Equation 10, we simulate flipping each bit in the plaintext coefficients. Let $j\in\{0,\dots,63\}$ denotes the bit position in a 64-bit unsigned integer. For each iteration, we flip the $j$-th bit of a coefficient, introducing an error with integer magnitude $e_{j}=2^{j}$. This represents the maximum jump in value induced by flipping bit $j$ in a coefficient, which is stored as an unsigned integer (which is the case of plaintext coefficients).
We repeat this process for all coefficients, grouping the results into $N=4$ blocks of 64 iterations each, and using $\Delta=2^{50}$. Let $\boldsymbol{z}^{\prime}$ be the decoded vector without errors, and $\boldsymbol{z}^{\prime\prime}$ the result after applying the bit flip and decoding. Using Equation 10, we express the error using the $L_{2}$ norm:
@@
where $\mathbf{e}_{i,j}$ is a zero vector of length $N$ with value $e_{j}$ at position $i$. We also take the real part of the vector before computing the $L_{2}$ norm, to simulate the encoding of real-valued data. This step is important not only to maintain consistency with the backend used in later sections, but also because it exposes a specific behavior that would otherwise not manifest when keeping the complex part.
Figure 4: $L_{2}$-norm of the error, measured as the difference between the golden (no bit-flip) output and the output obtained with a single bit-flip in the plaintext during encoding. The horizontal axis shows the bit-flip position. Parameters: $N=4$ and $\Delta=2^{50}$.
@@ III-C2 Effect of the Scaling Factor on Bit-Flip Error
Figure 5: $L_{2}$-norm of the error after a single bit-flip in the plaintext during encoding, for various scaling factors $\Delta$. The horizontal axis indicates the bit position $j$; the vertical axis is on a semi-log scale. Ring dimension: $N=4$.
As shown in Figure 5, increasing $\Delta$ significantly reduces the error magnitude and increases the number of resilient bits. This occurs because the impact of errors becomes noticeable only when the flipped bit corresponds to or exceeds the magnitude of $\Delta$. In other words, the larger the scaling factor, the more initial bits remain unaffected, yielding smaller error norms and higher bit-level fault tolerance.
This empirical observation is confirmed analytically by equation 11, which shows that the error is inversely proportional to the scaling factor. When an error $e_{i,j}$ is divided by $\Delta$, noise components with magnitudes smaller than the bit precision of $\Delta$ are effectively suppressed in the final result. This mathematical relationship explains why larger scaling factors provide enhanced fault resilience: they act as a natural filter that attenuates the impact of small-magnitude bit flips on the decoded output.
However, this resilience comes at the cost of reduced fractional precision in the encoded message (see Section II). The selection of $\Delta$ therefore requires careful consideration, balancing the competing demands of fault resilience and computational accuracy.
@@ III-D Theoretical Estimation of Error during Encryption using Vanilla CKKS
Under public-key encryption, a vector $\mathbf{m}$ is encoded into a polynomial $\boldsymbol{m}(X)$, encrypted as in Equation 3, and decrypted using Equation 4.
III-D1 Bit Flip in the Plaintext Before Encryption
Assuming that a bit flip introduces an error $\boldsymbol{e}_{i,j}$, as previously explained, the plaintext becomes:
$\boldsymbol{m}^{\prime}=\boldsymbol{m}+\boldsymbol{e}_{i,j}$ (12)
Encrypting $\boldsymbol{m}^{\prime}$ using the public key (Equation 3) yields:
$\begin{split}ct^{\prime}&=\bigl{(}[\,\boldsymbol{m}^{\prime}+p_{0}v+e_{1}\,]_{Q},\;[\,p_{1}v+e_{2}\,]_{Q}\bigr{)}\\ &=\bigl{(}[\,\boldsymbol{m}+\boldsymbol{e}_{i-1,j}+p_{0}v+e_{1}\,]_{Q},\;c_{1}\bigr{)}=\bigl{(}c_{0}^{\prime},\,c_{1}\bigr{)}.\end{split}$ (13)
The error affects only the ciphertext component $c_{0}^{\prime}$ in the coefficient $i$, with a maximum magnitude $|e|\leq Q/2$ due to modular arithmetic. Decryption (Equation 4) of such erroneous ciphertext yields:
$\boldsymbol{m}^{\prime\prime}=[\,c_{0}^{\prime}+c_{1}\,s\,]_{Q}=\boldsymbol{m}+\boldsymbol{e}_{i-1,j}+(e_{1}+e_{2}\,s)$ (14)
However, the error magnitude grows exponentially with the flipped bit position, as each bit represents a successive power of 2 in the binary representation. Consequently, an exhaustive bit-flip analysis across all coefficient positions, with L2 norm computation for each perturbation, yields the exponential error scaling pattern illustrated in Figure 4 (SectionIII-C). The bit-flip impact remains localized to the position $i$ and scales linearly with $e_{j}$.
III-D2 Bit Flip in the Ciphertext
If the bit flip occurs in the ciphertext, two sub-cases arise: